OpenStudy (anonymous):

\[3^{2-3x} = 4^{2x+1}\] Use common log and two decimal place approx.

OpenStudy (mani_jha):

(2-3x)log 3=(2x+1)log 4 Plug in the values of log 3 and log 4. Then it's just a linear equation to solve.

OpenStudy (anonymous):

I get: \[-x= \log 3^{-2} x 4 \div 3^{3}\div 4^{2} \]

OpenStudy (mani_jha):

\[2\log_{}3-3x \log_{}3=2x \log_{}4+\log_{}4 \] \[\log 3^{2}-\log4=2xlog 4+3xlog3\] \[\log_{} (9/4)=x(\log 16+\log 27)\] x=\[(\log(9/4)/\log(27/16)\] How did you get that? o.o

OpenStudy (anonymous):

Wasn't sure what to do with that "-x" when I pulled it out early on in the equation. Then I left in an extra x (before the 4.) So I'm confused about pulling out the x early on... do I keep that negative with the 3 log 3 - 2 log 4?

OpenStudy (mani_jha):

yes. don't keep it with x